Solve G(x) for the given domain.
G(x) = 3x 2 - 2x - 1
G(a + b) =

Answer by stanbon(75887)

Solve G(x) for the given domain.
G(x) = 3x^2 - 2x - 1
Domain: All Real Numbers.
-----

G(a + b) = 3(a+b)^2 -2(a+b) -1
----
= 3(a^2+2ab+b^2) -2a-2b -1
---
= 3a^2 + 6ab + 3b^2 -2a -2b -1
=================================
Cheers,
Stan H.